If an electron has an orbital angular momentum of 2.583E-34Js, what is the orbital quantum number for the state of the electron?
Answer #1

L=\sqrt{l\left ( l+1 \right )}\hbar\: \left ( 1 \right )

\hbar=\frac{h}{2\pi }

Extract l from (1): 6=l(l+1) --> l=2